OpenStudy (calculusxy):

MEDAL! PLEASE HELP! There is a discount of 40% for all golf club items. Rocco wants to use the rest of his money in his savings account to buy a set of golf irons that has an original price of $350. Ques: Rocco's friend works at the sporting goods store and gets 10% off the already discounted price. If Rocco's friend buys the golf irons for him, how much will his friend have to pay?

OpenStudy (calculusxy):

@jojob39 Does the 40% also count?

OpenStudy (calculusxy):

@Mertsj

OpenStudy (calculusxy):

@agent0smith

OpenStudy (calculusxy):

@e.mccormick

OpenStudy (mertsj):

350-40% of 350=350-140=210 Now his friend's discount: 10% of 210 = 21 So his friend pays 210-21=$189

OpenStudy (calculusxy):

So you do count the 40% right? And are golf club items part of golf irons? @Mertsj

OpenStudy (mertsj):

yes and yes

OpenStudy (e.mccormick):

The key is "10% off the already discounted price" Because it says already, you do the other discount first.

OpenStudy (calculusxy):

Thank you very much. I got that answer, but I was confused. Can you just help me with one more? Please?

OpenStudy (e.mccormick):

If it said an additional 10% off, you would work it differently.

OpenStudy (mertsj):

e.mccormick wants to help you.

OpenStudy (calculusxy):

I will definitely appreciate it :)

OpenStudy (e.mccormick):

I do? LOL

OpenStudy (calculusxy):

A sporting goods store offers a 40% discount on all golf clubs. Rocco spent 20% of the money in his saving's account on a golf putter. He paid $48 for the putter after the discount. B. Rocco wants to use the rest of his money in his savings account to buy a set of golf irons that has an original price of $350. Does Rocco have enough money to buy the golf irons?

OpenStudy (e.mccormick):

OK, so fist work out how much he has. If 20% was $48, how would you get back to the original value?

OpenStudy (calculusxy):

well there was this question on firsthand which was talking about how much was his balance before his purchase, i answered that as $240. so i think that using this information, i will be able to answer the question.

OpenStudy (e.mccormick):

Well, you can look at it right at the 20% of x = $48 too. That would work even if you were not given the value ahead of time. So with the original and what has been paid out, now you need to find how much he has at the moment.

OpenStudy (calculusxy):

i said that is was no. since i don't have the paper with me currently, i am guessing that i put the answer after the discounts as $210.

OpenStudy (e.mccormick):

Yes. $210 is more than he has left.

OpenStudy (calculusxy):

so it's no, i guess.

OpenStudy (e.mccormick):

You are correct.

OpenStudy (calculusxy):

thank you a lot. that's why you deserved the medal !!

OpenStudy (calculusxy):

..and fan behalf of me :)

OpenStudy (e.mccormick):

Well Mertsj has too many medals anyhow... =P

OpenStudy (calculusxy):

i guess. but anyways, you got it!!

OpenStudy (e.mccormick):

Well have fun!

OpenStudy (calculusxy):

u too :)